Diffusion Morning Report 02/24/06

Ted Berg tberg at snf.stanford.edu
Fri Feb 24 07:48:45 PST 2006


Good Morning all,
    Looking Ok this morning.HCL for epi is empty again-changing 
cylinders. BPSG is getting close to tube pull. HCl switched to A side.
Burn box cleaned this morning Everything else looks OK for now.

  	Date: 	2/24/2006 	THICKNESS-Change at
Tube Name/Number 	Status 	Comments 	 
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	Up 	  	 
Tylan 8-SiGe 	Up 	  	 18.2/30
Tylan 9-Poly 	Up 	  	2.25.15/30
Tylan 10-Nitride 	Up 	  	8.24/30
Tylan 11-BPSG-LTO 	Up 	  	26.16/30
Tylan 12-FGA 	Up 	  	 
  	  	  	 
Thermco 1 	Up 	  	 
Thermco 2 	Up 	  	 
Thermco Poly 	Up 	  	 
Thermco Nitride 	Down 	Testing 	 
  	  	  	 
Burn Box 	Up 	cleaned this morning 	 
ZYGO 	up 	  	  
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	down 	forming gas empty 	 
  	  	  	 
SiH4 for furnaces 	  	A on line 180 	 changeover working well 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20060224/7c609346/attachment.html>


More information about the diffusion mailing list